A Treasury bond futures contract has a settlement price of 89'08. What is the implied annual yield?

Price in whole units ?
Price in 32nds ?
Maturity in semiannual period ?
Semiannual coupon payments ?
Maturity value ?
Price ?
Semiannual implied yield ?
Annual implied yield ?
Answer #1

The answer is 7.01%

Calculations:

$1000*(89+8/32)/100 = 892.50

Value of contract = 892.50*100 = $89,250

In the financial calculator solve for I/YR. N = 40 (20 years of semi-annual payments).

PMT = 30, FV = 1000.

Thus PV = 892.50

Solving for I/YR = 3.504%

Then: 3.504%*2 = 7.01%

Thus implied annual yield = 7.01%